# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v2

@@ 5 Discussion
Despite MF-NPE’s advantages, the method comes with some challenges. First, the effectiveness of MF-NPE relies on the similarity between the low-fidelity and high-fidelity models. Fortunately, in many situations, domain experts will know beforehand whether low-fidelity models are poor approximations of high-fidelity models. Second, MF-NPE and MF-TSNPE inherit the limitations of NPE and TSNPE, respectively, in particular regarding the scalability of simulation-based inference to high-dimensional parameter spaces. How to balance exploration of high-dimensional parameter spaces and computational cost in a simulation-based inference setting remains a topic of active research.
We identify two promising research directions in the realm of multifidelity simulation-based inference. First, we expect that the scalability and expressivity of MF-NPE can be improved by utilizing the same approaches of multifidelity and transfer learning presented here with neural density estimators other than normalizing flows, such as diffusion models (Gloeckler et al., 2024). Second, similar to past efforts in developing a benchmark for simulation-based inference, it will be beneficial for the SBI community to develop a benchmark for multifidelity problems, with new tasks, algorithms and evaluation metrics. This will promote rigorous and reproducible research and catalize new developments in multifidelity SBI, and in SBI more generally. Our work and codebase are a step in this direction.
@@ Conclusion
Overall, MF-NPE is a method for simulation-based inference that leverages low-fidelity models and transfer learning to infer the parameters of costly high-fidelity models, thus providing an effective balance between computational cost and inference accuracy.
Acknowledgements
@@
We thank Karthik Sama, Najlaa Mohamed, Guy Moss, Marcel Nonnenmacher and Pierre Vanvolsem for discussions. Anastasia N. Krouglova was supported by an FWO grant (G097022N). Hayden R. Johnson was supported by an FWO grant (G053624N). Basile Confavreux was supported by a Schmidt Science Polymath Award to Andrew Saxe, the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3850). Michael Deistler was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC 2064 – Project number 390727645), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039A) and the European Union (ERC, “DeepCoMechTome”, ref. 101089288). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Michael Deistler is a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Further experimental details
A.1 Training procedure
@@
All methods and evaluations were implemented in PyTorch (Paszke et al., 2019). We used Zuko package (Roset, 2024) implementations of Neural Spline flows (NSF) (Durkan et al., 2019) and SBI package (Boelts et al., 2024) for additional functions. The parameters were logit-transformed for numerical stability, and the summary statistics were z-scored to improve the performance of the Normalizing Flows.
The Normalizing Flows had 5 transformations, each parametrized with 50 hidden units and 8 bins, the default settings in the SBI package (Boelts et al., 2024). All results were obtained over an ensemble of 10 networks with a fixed training dataset and a random train-test split.
A.2 Data generation and transformations for increased network performance
Appendix B Task 1: Ornstein-Uhlenbeck process
B.1 High-fidelity model: Ornstein-Uhlenbeck Process
@@
Figure 6: The four parameters of the Ornstein-Uhlenbeck process: the mean $\mu$, standard deviation $\sigma$, convergence rate $\gamma$, and $\mu_{\mathrm{offset}}$, which is the difference between the initial condition $X(0)$ and the mean.
@@
$f_{\text{exact hi}}(\boldsymbol{X}\mid\mu,\sigma,\gamma,\mu_{\text{offset}})=\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{(x-(\mu+\mu_{\text{offset}}))^{2}}{2}\right\}\prod_{t=1}^{n}\frac{1}{\sqrt{\pi g}\sigma}\exp\left\{-\frac{1}{g\sigma^{2}}\left(\left(\mu-X_{t}\right)-\sqrt{1-\gamma g}\left(\mu-X_{t-1}\right)\right)^{2}\right\}$
@@
B.2 Low-fidelity model: i.i.d. Gaussian Samples
At convergence, the distribution over $X_{t}$ approaches a Gaussian distribution with mean $\mu$ and standard deviation $\frac{\sigma}{\sqrt{2\gamma}}$. In our setup, we chose a low-fidelity model that corresponds to time-independent random draws from a Gaussian distribution with mean $\mu_{\textrm{lo}}$ and standard deviation $\sigma_{\textrm{lo}}$:
@@
The summary statistics from both the high-fidelity and low-fidelity models consisted of 10 subsamples drawn from a trace of 100 timesteps. Parameters and summary statistics are described in Fig. 7.
B.3 Posterior distribution over different dimensionalities
When the low-fidelity has lower number of parameters than the high-fidelity model, we pre-train the low-fidelity density estimator with dummy parameter values in the dimensions exclusive to the high-fidelity model. Effectively, the pre-conditioning with the dummy variables leads to the pre-trained neural density estimator to effectively estimate the prior distribution over the respective parameters (Figure 8A). Despite the different dimensionalities, the posterior estimates from MF-NPE are closer to the ground-truth posterior than NPE (Figure 8B,C).
[…]
Appendix C Task 2: Multicompartmental single neuron model

### v3

@@ 5 Discussion
Limitations
Despite MF-NPE’s advantages, the method comes with some challenges. First, the effectiveness of MF-NPE relies on the similarity between the low-fidelity and high-fidelity models. Fortunately, in many situations, domain experts will know beforehand whether low-fidelity models are poor approximations of high-fidelity models. Second, MF-NPE and MF-TSNPE inherit the limitations of NPE and TSNPE, respectively, in particular regarding the scalability of simulation-based inference to high-dimensional parameter spaces. How to balance exploration of high-dimensional parameter spaces and computational cost in a simulation-based inference setting remains a topic of active research. Third, A-MF-TSNPE requires the training of an ensemble of density estimators, which leads to substantial computational costs in training and hyperparameter tuning. This method should therefore only be preferred in cases where the cost incurred in simulations outweighs the training cost. We estimate this to be the case for the tasks with the multicompartment neuron model and the spiking network model, for which the cost of one simulation and the training of one density estimator are comparable in certain settings (e.g., on the order of minutes, for a network trained on $10^{3}$ samples).
Future work
We identify three promising research directions for multifidelity simulation-based inference. First, we expect the scalability and expressivity of MF-NPE could be improved by utilizing the same approaches of multifidelity and transfer learning presented here with neural density estimators other than normalizing flows, such as diffusion models (Gloeckler et al., 2024). Second, we assumed a negligible cost for low-fidelity simulations, and future work should address how to optimally allocate low- and high-fidelity simulations under a fixed computational budget. Third, similar to past efforts in developing a benchmark for simulation-based inference, it will be beneficial for the SBI community to develop a benchmark for multifidelity problems, with new tasks, algorithms and evaluation metrics. This will promote rigorous and reproducible research and catalyze new developments in multifidelity SBI, and in SBI more generally. Our work and codebase are a step in this direction.
@@ Conclusion
Overall, MF-(TS)NPE is a method for simulation-based inference that leverages low-fidelity models and transfer learning to infer the parameters of costly high-fidelity models, thus providing an effective balance between computational cost and inference accuracy.
6 Reproducibility statement
The training and simulation costs for all tasks and SBI methods, as well as a detailed description of the experimental setup, are described in Appendices D.1 and K. The corresponding code and data will be publicly released upon publication.
Acknowledgments
@@
We thank Karthik Sama, Najlaa Mohamed, Guy Moss, Marcel Nonnenmacher and Pierre Vanvolsem for discussions. We also thank the staff of the VIB Data Core for their support. Anastasia N. Krouglova was supported by an FWO grant (G097022N). Hayden R. Johnson was supported by an FWO grant (G053624N). Basile Confavreux was supported by a Schmidt Science Polymath Award to Andrew Saxe, the Sainsbury Wellcome Centre Core Grant from Wellcome (219627/Z/19/Z) and the Gatsby Charitable Foundation (GAT3850). Michael Deistler was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC 2064 – Project number 390727645), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039A) and the European Union (ERC, “DeepCoMechTome”, ref. 101089288). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Michael Deistler is a member of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix A Usage of LLMs
LLM usage was minimal, limited to grammar refinement, sentence shortening, code cleanup and discovering papers outside our main domain.
Appendix B When does pre-training help?
In this section, we empirically investigate conditions in which low-fidelity pre-training facilitates effective transfer learning in approximating the high-fidelity posterior. Concretely, we study MF-NPE in the special case where the low- and high-fidelity models describe the same stochastic process but differ through a controlled perturbation. This setting lets us vary the amount of shared information between the two simulators in a principled way and quantify the effect on downstream performance.
We hypothesize that the effectiveness of pre-training is associated with two primary factors:
  1. 1.
    Mutual information between the low- and high-fidelity simulators.
  2. 2.
    Representational coherence, i.e., similarity in how task-relevant information is encoded.
To test this hypothesis, we consider the OU2 process under additive Gaussian noise and/or an information-preserving shift in data representation.
As a shift in data representation, we consider a reversal of the coordinate order of the OU-process output. When $x\in\mathbb{R}^{D}$, we can define this using an anti-diagonal permutation matrix
$T(x)=J\,x,\qquad J_{ij}=\begin{cases}1,&j=D-i+1,\\ 0,&\text{otherwise}.\end{cases}$
which simply reverses the coordinate ordering of $x$. Notably, this transformation is invertible and thus information-preserving.
B.1 Controlled experiment: perturbed OU2 process
In our experiment, we define the following generative process:
  1. 1.
    Sample simulation parameters from the prior: $\mu,\sigma\sim p(\mu,\sigma).$
  2. 2.
    Generate a ”high-fidelity” observation given the sampled parameters: $X_{\text{H}}\sim p\!\left(X\,\middle|\,\mu,\sigma\right)$
  3. 3.
    Draw a noise term: $\varepsilon\sim\mathcal{N}\!\left(0,\,\delta^{2}\right)$
  4. 4.
    Generate a ”low-fidelity” observation: $X_{\text{L}}\sim p\!\left(X\,\middle|\,\mu,\sigma+\varepsilon\right)$
  5. 5.
    Apply the transform: $X_{\text{L}}^{\mathrm{inv}}\;=\;T\!\bigl(X_{\text{L}}\bigr)$, where $T$ is the inversion defined above.
As mutual information is invariant to invertible transformations, we can write
$\mathbb{I}[X_{H};X^{\text{inv}}_{L}]=\mathbb{I}[X_{H};X_{L}]=\mathbb{H}[X_{H}]+\mathbb{H}[X_{L}]-\mathbb{H}[X_{H},X_{L}]\\$ (3)
The quantity $\mathbb{I}[X_{H};X_{L}]$ decreases monotonically with the noise scale $\delta$. At the same time, applying the inversion leaves the information content unchanged while disrupting representational coherence. Fig. 4 illustrates how each manipulation affects the uncertainty coefficient
$\mathbb{UC}[X_{H}|X_{L}]=\frac{\mathbb{I}[X_{H};X_{L}]}{\mathbb{H}[X_{H}]},$
which we estimate empirically using MINE (Belghazi et al., 2018) (Fig. 4B), and MF-NPE performance under a fixed simulation budget of $10^{4}$ low-fidelity simulations and $10^{2}$ high-fidelity simulations (Fig. 4C).
Figure 4: (A) Schematic figure representing lower bound on transfer error ($1/\text{MF-NPE{} performance}$) as a function of mutual information between the low- and high-fidelity models, given a fixed simulation budget. (B) Uncertainty coefficient monotonically decreases with noise parameter $\delta$ and is invariant to data inversion. (C) Empirical results with MF-NPE support the hypothesis that transfer performance is dependent on both mutual information and representational coherence. Note that NPE—with the same high-fidelity simulation budget of $10^{2}$—has similar performance as MF-NPE in the case where the low- and high-fidelity models have low mutual information.
In agreement with our hypothesis, our results suggest the effectiveness of MF-NPE is associated with both the mutual information and representational coherence between low- and high-fidelity simulators (Fig. 4C). Specifically, we find that mutual information is necessary for effective transfer learning, but is not sufficient. Perturbations that preserve information (e.g., invertible transformations) can also critically affect transfer learning performance. Thus, effective pre-training strategies should prioritize high mutual information and coherent representation to maximize downstream performance.
Appendix C Proof of convergence of the NPE log-likelihood loss
Let $\theta_{i}\sim p(\theta_{i})$ be samples from the prior of a high-fidelity model, and $x_{i}\sim p(x|\theta_{i})$ be the respective high-fidelity simulations. In NPE, we define the loss function as the negative log likelihood:
$\mathcal{L}(\phi)=-\frac{1}{N}\sum_{i}^{N}\log q_{\phi}(\theta_{i}|x_{i}),$ (4)
where $\theta_{i}$ are samples from the prior distribution, $x_{i}$ are the respective simulations (i.e., samples from $p(x|\theta_{i})$), and $\boldsymbol{\phi}$ are the parameters of the neural density estimator to be optimized. If we let the number of samples $\theta_{i}$ (and respective simulations) $N\rightarrow\infty$:
[…]
where $C$ is a constant with respect to $\phi$. Minimizing $\mathcal{L}(\phi)$ with respect to $\phi$ is thus equivalent to minimizing the KL divergence between the true posterior distribution and the estimated posterior in the limit of an infinite number of high-fidelity samples.
Appendix D Further experimental details
D.1 Training procedure
@@
All methods and evaluations were implemented in PyTorch (Paszke et al., 2019). We used the Zuko package (version 1.4.0, MIT License) [^1] https://github.com/probabilists/zuko(Roset, 2024) to implement the normalizing flow, based on the Neural Spline Flows (NSF) architecture (Durkan et al., 2019), and the SBI package (version 0.24.0, Apache 2.0 license) [^2] https://github.com/sbi-dev/sbi (Boelts et al., 2024) for additional functions. The parameters used to generate simulations were logit-transformed for numerical stability, and the summary statistics were z-scored to improve the performance of the normalizing flows. The loss function is the negative-log likelihood, and the optimization function is the Adam optimizer (Kingma & Ba, 2017).
The Neural Spline Flow (NSF) architecture consists of 5 transformations, each parametrized with 50 hidden units and 8 bins. The batch size was set to 200, and the learning rate to $5\times 10^{-4}$. The train-validation fraction is $0.1$, and training of the NSF utilized an early stopping criterion with a patience of 20 epochs for the early stopping criterion. The settings described above are all default settings of the SBI package at the time of the method’s development (Boelts et al., 2024).
Note, the stopping criterion follows the default configuration of the SBI package, which is defined as follows: Let $E$ be the error function of the training algorithm (negative log likelihood), $E_{val}(t)$ the validation error at epoch $t$, which is used by the stopping criterion. The value $E_{opt}(t)$ is the lowest validation set error obtained in epochs up to $t$:
$E_{opt}(t):=\min_{t^{\prime}\leq t}E_{val}(t^{\prime})$ (6)
The early stopping criterion $S$ terminates training once the validation error $E_{\text{val}}(t)$ has increased for $p$ consecutive epochs (the patience parameter). At this point, the model corresponding to the lowest validation error observed that far, $E_{\text{opt}}(t)$, is selected and returned.
Rather than fixing the number of training epochs, the idea behind early stopping is that when the validation error has increased not only once, but over $p$ consecutive steps, such an increase indicates a stage of overfitting (Prechelt, 2002). Note that if the patience is too small, underfitting might occur, and training may terminate too early due to stochastic fluctuations in the loss. Similarly, overfitting might likely occur when the patience is set to excessively high numbers (especially with a low number of simulations, since the loss function is typically more variable in this setting).
For the fine-tuning step of MF-NPE, no network weights were frozen. This choice has been purposely made to maintain full flexibility of the network to adapt to the high-fidelity model.
For the evaluation of A-MF-TSNPE, we used 5 rounds of active sampling, where $80\%$ of the high-fidelity dataset was used for standard MF-NPE training, and $20\%$ was split across the rounds of active sampling. The active samples were selected using the acquisition function over an ensemble of 5 networks.
For a fair performance comparison, all methods were trained on the same datasets and evaluated on the same observations $x_{o}$. All amortized results were obtained over 10 network initializations, and all non-amortized results over 1 or 10 network initializations (depending on the computational cost of the task). We evaluated the methods over 30 observations for the C2ST metric, more than the 10 observations chosen previously for benchmarking (Lueckmann et al., 2021). This choice is motivated by our focus on evaluating the methods in low-data regimes, where greater certainty is required. The performance on the L5PC neuron task was evaluated with the metric NLTP and over 100 $x_{o}$’s. Here, the performance of the amortized methods was averaged over 10 network initializations, and in the non-amortized methods over 1 network initialization, since training had to be performed for each individual $x_{o}$. The performance of the methods on the recurrent spiking network task was averaged over 10 network initializations and evaluated over 262,008 observations, which was the maximum number of available samples for this high-dimensional problem.
Appendix E Tasks
E.1 OU process
The Ornstein-Uhlenbeck (OU) process is a high-fidelity model with 2 to 4 free parameters that contains a temporal structure in the observations. As a low-fidelity model, we chose i.i.d. samples from a Gaussian distribution (unstructured vector), parametrized by the mean and standard deviation. This setting makes it well-suited to examine the impact of parameter space overlap between the low- and high-fidelity models, as well as the impact of a systematic bias in the posterior of the low-fidelity model on transfer learning.
@@ E.1 OU process
Figure 5: The four parameters of the Ornstein-Uhlenbeck process: the mean $\mu$, standard deviation $\sigma$, convergence rate $\gamma$, and $\mu_{\mathrm{offset}}$, which is the difference between the initial condition $X(0)$ and mean $\mu$.
High-fidelity model
@@
$f_{\text{exact hi}}(\boldsymbol{X}\mid\mu,\sigma,\gamma,\mu_{\text{offset}})=\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{(x-(\mu+\mu_{\text{offset}}))^{2}}{2}\right\}f_{\text{exact hi}}(\boldsymbol{X}\mid\mu,\gamma,\sigma)$
@@
Low-fidelity model
As a low-fidelity model, we use i.i.d. Gaussian Samples. At convergence, the distribution over $X_{t}$ approaches a Gaussian distribution with mean $\mu$ and standard deviation $\frac{\sigma}{\sqrt{2\gamma}}$. In our setup, we chose a low-fidelity model that corresponds to time-independent random draws from a Gaussian distribution with mean $\mu_{\textrm{lo}}$ and standard deviation $\sigma_{\textrm{lo}}$:
@@
Prior $\mu\sim\mathcal{U}(0.1,3),\quad\sigma\sim\mathcal{U}(0.1,0.6),\quad\gamma\sim\mathcal{U}(0.1,1),\quad\mu_{\text{offset}}\sim\mathcal{U}(0,4)$
HF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{101}),\quad x_{0}\sim\mathcal{N}(\mu+\mu_{\text{offset}},1),$ where
$dx_{t}=\gamma(\mu-x_{t})dt+\sigma dW_{t}$
LF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{10}),\quad x_{i}\sim\mathcal{N}(\mu_{\text{lo}},\sigma_{\text{lo}}^{2}),$
HF Dimensionality $\theta\in\mathbb{R}^{2-4},\quad\mathbf{x}\in\mathbb{R}^{101}$, $\quad U(\mathbf{x})\in\mathbb{R}^{10}$
LF Dimensionality $\theta\in\mathbb{R}^{2}$, $\quad\mathbf{x}\in\mathbb{R}^{10},\quad U(\mathbf{x})\in\mathbb{R}^{10}$
References (Holý & Tomanová, 2022; Carter & Strey, 2023; Kou et al., 2012)
For the two-dimensional experiment, the free parameters $\gamma,\mu_{\text{offset}}$ have been fixed to $\gamma=0.5$ and $\mu_{\text{offset}}=3.0$. For the three-dimensional-experiment, only $\mu_{\text{offset}}=3.0$. The summary statistics $U(x)$ from the high-fidelity model consists of 10 uniformly distributed subsamples drawn from a trace of 101 timesteps. Parameters and summary statistics are illustrated in Figures 5 and 6.
E.2 Posterior distributions over OU process
Figure 7: Posterior density estimates for a single observation from the OU process with two free parameters (OU2). The orange contour lines contain 68% of the probability mass of the true posterior distribution.
E.3 OU process with varying parameter space
We present a comparison of our multifidelity approaches to NPE and MF-ABC, with different numbers of pre-trained low-fidelity simulations. MF-NPE3 is pre-trained on a low-fidelity dataset of size $10^{3}$, while MF-NPE4 and MF-NPE5 use datasets of $10^{4}$ and $10^{5}$ low-fidelity simulations, respectively. The MF-ABC results suggest that neural density approaches scale better to complex problems (Frazier et al., 2024).
Figure 8: MF-NPE benefits from larger low-fidelity datasets. We ran MF-ABC with hyperparameters $\boldsymbol{\epsilon}=(1,1)$ and $\boldsymbol{\eta}=(0.9,0.3)$ (more details in Appendix F.1.1). All variants of our method perform better than MF-ABC and NPE.
E.4 SLCP
Simple Likelihood Complex Posterior (SLCP) is s a benchmark inference task that has been artificially designed to have a simple likelihood, but a very non-trivial 5-dimensional posterior to infer. In this example, we study the impact of multifidelity in cases where the dimensionality of the parameter space differs between the low-fidelity and high-fidelity models.
High-fidelity model
The SLCP problem involves five parameters. The prior distribution is uniform across a five-dimensional parameter space, and the observations consist of four two-dimensional samples drawn from a Gaussian distribution. Both the mean and the variance of this Gaussian depend on the parameters through nonlinear mappings. The high-fidelity model follows the code in the SBI benchmarking paper (Lueckmann et al., 2021).
Low-fidelity model
In the low-fidelity model, we experimented with the effect of different numbers of parameters on the inference quality. We fixed $m_{\theta}=0$, and kept the parameters of $S_{\theta}$ free.
Prior $\mathcal{U}(-3,3)$
HF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{4}),\quad x_{i}\sim\mathcal{N}(\mathbf{m}_{\theta},\mathbf{S}_{\theta}),$
where $\mathbf{m}_{\theta}=\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},\quad\mathbf{S}_{\theta}=\begin{bmatrix}s_{1}^{2}&\rho s_{1}s_{2}\\ \rho s_{1}s_{2}&s_{2}^{2}\end{bmatrix}$,
with $s_{1}=\theta_{3}^{2},\;s_{2}=\theta_{4}^{2},\;\rho=\tanh(\theta_{5}).$
LF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{4}),\quad x_{i}\sim\mathcal{N}(0,\mathbf{S}_{\theta}),$
where $\mathbf{S}_{\theta}=\begin{bmatrix}s_{1}^{2}&\rho s_{1}s_{2}\\ \rho s_{1}s_{2}&s_{2}^{2}\end{bmatrix}$,
with $s_{1}=\theta_{3}^{2},\;s_{2}=\theta_{4}^{2},\;\rho=\tanh(\theta_{5}).$
HF Dimensionality $\theta\in\mathbb{R}^{5},\quad\mathbf{x}\in\mathbb{R}^{8}$
LF Dimensionality $\theta\in\mathbb{R}^{3},\quad\mathbf{x}\in\mathbb{R}^{8}$
References (Papamakarios et al., 2019; Hermans et al., 2020)
(Durkan et al., 2020; Greenberg et al., 2019; Lueckmann et al., 2021)
(Thiele et al., 2025)
E.5 SIR
The Susceptible, Infected, and Recovered (SIR) model is a classical epidemiological benchmark example that captures the spread of infectious diseases through three interacting compartments: Susceptible (S), Infectious (I), and Recovered (R). Its dynamics are governed by the system of ordinary differential equations. The model is parameterized by two rates: the infection rate $\beta$ and the recovery rate $\gamma$. We investigate how multifidelity addresses the partly observed dynamics of the model. Rather than observing the three dynamics of the SIR model (following the setup of the SBI benchmarking (Lueckmann et al., 2021), we assume that no dynamics regarding the recovered subjects are known (SI model).
Low-fidelity model
In the low-fidelity model, we assume no information is available about the dynamics of recovered individuals. The total population size and the initial conditions are kept consistent with the high-fidelity model.
Bounded domain $0.001,3]^{2}$
Prior $\beta\sim\text{LogNormal}(\log(0.4),0.5),\quad\gamma\sim\text{LogNormal}(\log(0.125),0.2)$
HF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{50}),\quad x_{i}=I_{i}/N\text{ equally spaced,}$
II is simulated from $\dfrac{dS}{dt}=-\beta\dfrac{SI}{N},\quad\dfrac{dI}{dt}=\beta\dfrac{SI}{N}-\gamma I,\quad\dfrac{dR}{dt}=\gamma I$
LF Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{50}),\quad x_{i}=I_{i}/N\text{ equally spaced,}$
II is simulated from $\dfrac{dS}{dt}=-\beta\dfrac{SI}{N},\quad\dfrac{dI}{dt}=\beta\dfrac{SI}{N}-\gamma I,$
Dimensionality $\theta\in\mathbb{R}^{2},\;x\in\mathbb{R}^{3\times 161},\;U(\mathbf{x})\in\mathbb{R}^{10}$
Fixed parameters Population size $N=10^{6}$, duration of task $T=160$ days.
Initial conditions: (S(0),I(0),R(0))=(N1,1,0)(S(0),I(0),R(0))=(N-1,1,0)
References (Lueckmann et al., 2021; Greenberg et al., 2019)
(Hermans et al., 2020; Durkan et al., 2020)
Summary statistics $U(x)$ are 10 subsamples from the $I$ trace.
E.6 Image example
We apply our method to a problem with high-dimensional observations, and explore the benefits of transfer learning in combination with embedding networks. The high-fidelity model is a 256x256 image, while the low-fidelity model has a resolution of 32x32. An example of both simulator outputs is shown in Fig. 9.
High-fidelity model
The Gaussian Blob image example contains high-dimensional observations that have been embedded with a CNN embedding from the SBI package (Boelts et al., 2024). The model renders a 2D image, which we modeled as a 256 x 256 pixel image of a Gaussian blob, and aiming to infer three parameters ($\mu_{\text{off}},\sigma_{\text{off}},\gamma$): the horizontal and vertical displacements of the blob, and its contrast (Lueckmann et al., 2019). The image is a grey-scale and is generated through a binomial distribution with a total count of 255 and probability $p_{ij}$, as described in Lueckmann et al. (2019).
Low-fidelity model
In our setup, the low-fidelity model generates a spatially low-resolution dataset (32x32 image). We upscale these images using interpolation techniques and provide the resulting low-resolution inputs to the embedding network $U(x)$.
Prior HF $x_{\text{off}},y_{\text{off}}\sim\mathcal{U}(0,256),\quad\gamma\sim\mathcal{U}(0.2,2)$
Prior LF $x_{\text{off}},y_{\text{off}}\sim\mathcal{U}(0,32),\quad\gamma\sim\mathcal{U}(0.2,2)$
Simulator $\mathbf{x}|\theta=(x_{1},\ldots,x_{1024}),\quad$ where,
$I_{xy}\sim\text{Bin}(\cdot|255,p_{xy})$
$p_{xy}=0.9-0.8\exp^{-0.5(r_{xy}/\sigma^{2})^{\gamma}}$
$r_{xy}=(x-x_{\text{off}})^{2}+(y-y_{\text{off}})^{2}$
Dimensionality HF $\theta\in\mathbb{R}^{3},\;x\in\mathbb{R}^{256\times 256}$, $U(x)\in\mathbb{R}^{32}$
Dimensionality LF $\theta\in\mathbb{R}^{3},\;x\in\mathbb{R}^{32\times 32}$, $U(x)\in\mathbb{R}^{32}$
Fixed parameters Standard deviation $\sigma_{\text{lf}}=2$, $\sigma_{\text{hf}}=12$
References (Lueckmann et al., 2019)
Appendix F Gaussian Blob evaluation
Figure 9: Five examples of generated images with the Gaussian Blob across the two fidelities, with (a) the original 256x256 high-fidelity simulations, (b) the upsampled 32x32 low-fidelity simulations.
Figure 10: Method comparison with NLTP and NRMSE for the Gaussian Blob task. Evaluated over 10000 observations.
F.1 Data generation and transformations for increased network performance
F.1.1 Multifidelity Approximate Bayesian Computation (MF-ABC)
We translated into Python a publicly available Julia implementation of the multifidelity ABC algorithm (Prescott & Baker, 2020). In our setup, the adaptive sampling scheme of MF-ABC selected approximately $30\%$ of the batch size as high-fidelity samples in the OU2 and OU3 tasks, and $50\%$ in the OU4 task. To ensure consistency with our neural network experiments, we z-scored the simulator output before inference. We also explored the effect of varying the acceptance threshold $\epsilon$. We found that the hyperparameters slightly affect the performance of MF-ABC, but that MF-NPE always shows superior performance than MF-ABC (Figure 11). However, MF-ABC has several other hyperparameters to tune. We cannot exclude the hypothesis that larger performance gains could be obtained from such an approach by a more extensive hyperparameter search.
Figure 11: C2ST results for MF-ABC with varying hyperparameters $\boldsymbol{\epsilon}$. Mean and $95\%$ confidence interval.
MF-ABC posteriors
ABC-based methods typically require a significantly larger number of samples for convergence (Lueckmann et al., 2021; Frazier et al., 2024). In line with previous studies, we find that $10^{4}$ samples are not yet enough for MF-ABC to converge to a good estimate of the posterior in the OU2 task.
Figure 12: Comparison between MF-ABC posterior estimates and the true posterior. Results for the Ornstein-Uhlenbeck process with two free parameters. Posterior estimates are shown for varying numbers of high-fidelity simulations (50, 100, $10^{3}$, and $10^{4}$).
Appendix G Task 2: Multicompartmental single neuron model